Complete the shear moment diagrams.

600 N 400 N/m С PA B -2 m -2 m 2 m
Answer #1

bon FBD- Given Beam is present at point is internal hinge So brea brewning into to parts and drawing the FBD for the ReactioRc UVON this from equation RA 1400 - Hoo RA = 1000 N B Ffr/ Set Paling moment foom left side about point Em3=0 =) -MATRAX 4 G400-400 (9-4)=0 40019-4) = 400 71-4 = 1 725m + 400 (SFDB (SF) = +400-400 (6-4) -400N + 100E N +1 OOON TULON tuto N + Hiem oltEn Sedon B to C Mc=0 MB=0 2 Mx 2800 + 1000 x x - Gero (2-2) -400 (7-4 2 { from left side Finding the point where it is zero.+ 2050 Nm 20 am Im im 3 C -800nm 2800 Nm

Please forgive me for any calculation mistake and do comment for any short of queries.

Please give thumbs up if you liked my efforts.